A bullet fired in to a target loses half of its velocity after penetrating 25cm.how much further will it penetrate before coming to rest?
1
Expert's answer
2021-08-19T14:21:36-0400

We know that

"v^2=u^2+2as"

"\frac{u^2}{4}=u^2+2a\times25\\-\frac{3u^2}{4}=50a\\a=-\frac{3}{200}u^2"

"u^2=v^2+2as"

Put a value

u=0

"0=u^2-2\times\frac{3}{200}u^2s\\s=\frac{100}{3}cm"